# Field equation of thermodynamic gravity and galactic rotational curves

### v2

@@ 3.3. Discussion
NGC 3198 is a barred spiral galaxy known for its prototypical flat rotation curve, [32, 33, 34]. It was chosen because all three investigated theories provide a good fit for the observed velocity data. Additional galaxies will be investigated in future work. It is well-studied in the context of MOND, as the standard value of $a_{0}$ cannot predict the rotational velocity curve unless the distance of the galaxy is adjusted. Randriamampandry and Carignan found an acceptable MOND fit with $a_{0}$ about half the standard value. A much smaller distance is needed than the Cepheid-distance of $13.8$ Mpc [30]. As seen in Fig. 8, thermodynamic gravity can very similarly match the rotation curves of the ISO DM model and also of MOND, with $a_{0}$ being allowed to vary. A notable difference that TG seems to be more sensitive to the density distribution of the barionic matter, as it is apparent in the region where the density of the atomic hydrogen is negative, indicating the oversimplified nature of the spherical symmetric mass modell.
@@
In Fig. 9 one can see separately the contribution of the neutral hydrogen to the velocity curve. One can see, that the contribution of atomic hydrogen is necessary for a proper modelling of the flat part.
@@ 3.4. Mass estimation
The datasets in the THINGS sample provide masses for the fitted mass models and using the fitted ISO DM parameters, the mass of the ISO DM halo can be calculated by integrating it up to the edge of the galaxy. For the purposes of this estimation, the edge of the galaxy will be considered to be at the last observational velocity curve datapoint ($r_{d}$). Using the insights from eqs. 13 and 16, by considering the fitted K parameters for the galaxy, it is possible to estimate the apparent asymptotic mass of the galaxy by setting the acceleration inferred from the edge of the galaxy to equal to the vacuum solution:
@@
In the case of $K=0$, eq. 56 should approximately yield the gravitating mass required for the observed velocity for the galaxy and is expected from the total mass inside $r_{d}$ with DM included. Therefore, we can call $M_{aa,K=0}$ as apparent total asymptotic mass of the galaxy. The values including dark matter or $M_{aa}$ are rounded to integers due to the large uncertainties included. Therefore, $M_{BM+ISO}$ inherits the uncertainty of $M_{ISO}$. The results are shown in Table 1:
Galaxy $M_{SD}$ $M_{HI}$ $M_{total}$ $M_{ISO}$ $M_{BM+ISO}$ $M_{aa,K=0}$ $M_{aa}$
NGC 3198 $31.01$ $15.37$ $46.38$ $136\pm 19$ $183\pm 19$ $196$ $110$
Table 1. The mass components and estimates for ISO DM mass, all values are in the units of $10^{9}M_{\odot}$. $M_{SD}$ is the mass of the stellar components, $M_{HI}$ is the mass of neutral hydrogen, $M_{total}$ is the total mass of $M_{SD}$ and $M_{HI}$, $M_{ISO}$ is the mass of the ISO DM halo, integrated to $r_{d}$, $M_{BM+ISO}$ is the total mass, including baryonic and DM components, and $M_{aa}$ is the apparent asymptotic mass from the thermodynamic theory.
We can see, that the total baryonic and dark matter mass of NGC 3198, $M_{BM+ISO}$, is about the same as the apparent total mass of TG (expected at large distances), but the apparent asymptotic mass with the galactic $K$ value, $M_{aa}$, while differing from the ISO DM mass, up to $r_{d}$ results in the same gravitational effect. The characteristic $R$ spatial distance of the crossover, as pictured on Fig. 4 and detailed in [20] can be obtained as
(57) $R=KGM_{aa},$
@@
Using the $M_{aa,K=0}=196\cdot 10^{9}M_{\odot}$ mass for this scale, the expected characteristic crossover distance of this galaxy is $R\approx 60$ kpc, outside even of the extended data points of [35]. That is the distance where the flat velocity curve is expected to start declining and tending to a Newtonian one, which is determined by the vacuum solution of the Poisson equation with the apparent total asymptotic mass.
@@ 3.5. Conclusions
In this work, the analytical solutions of the nonlinear stationary field equation of TG were presented for the gravitational field in the cases of vacuum and constant matter density. Using the staggered grids-type discretisation, a numerical relaxation method was developed to solve the dissipative field equation of the theory. The stationary solution was obtained in the limit of the relaxation. The rotational velocity curve was calculated for galaxy NGC 3198 from the THINGS dataset, with only the baryonic matter contributing to the matter density source term [29]. It is remarkable, that NGC 3198 was previously investigated for the possible fits by MOND [30], and the ideally universal constant $a_{0}$ proved inadequate, as the best fits were only obtained when this parameter was allowed to vary, or the established distance of the galaxy was revised. The $K$ parameter in the theory investigated by this work arises from the coupling between the gravitational and mechanical thermodynamic forces and fluxes and, thus, naturally varies across the galaxies, depending on their smaller-scale dynamics and characteristics. Based on the developed numerical method, a next step is to extend the analysis to other galaxies of the THINGS survey, with a detailed comparison of the alternative models. It is also natural to investigate thermodynamic gravity on the scale of galaxy clusters. It may be the case that on much larger scales, the theory has effects similar to Newtonian or general relativistic gravity, except the apparent mass is larger due to the thermodynamic coupling, with a crossover between the regimes [20].
@@
[…]
@@ Acknowledgements
We would like to thank Professor W. J. G. de Blok for providing us with the THINGS data [29]. We would also like to thank Professor S. Abe for providing insightful remarks regarding the theory.
@@
The data for the galaxy NGC 3198 used in this paper is based on the work of de Blok [29], DOI.: 10.1088/0004-6256/136/6/2648.

### v3

@@ 3.3. Discussion
NGC 3198 is a barred spiral galaxy known for its prototypical flat rotation curve, [40, 41, 42]. It was chosen because all three investigated theories provide a good fit for the observed velocity data. It is remarkable that first MOND based rotation curve fits give acceptable result only with varying MOND acceleration parameters [43, 39]. However, a more sophisticated modelling methodology considers the observation uncertainties and results in a good modelling also with the fixed, universal $a_{0}$, as in the MOND EFE model of Li et al. [35, 36]. In our work, thermodynamic gravity can very similarly match the rotation curves as the DC14 model DM and also as MOND EFE.
@@ 3.4. Mass estimation
The DC14 model total halo mass can be estimated as:
(62) $M(r)=4\pi r_{s}^{3}\rho_{s}\frac{1}{\alpha}\left[B(a,b+1,\epsilon)+B(a+1,b,\epsilon)\right].$
For the purposes of this estimation, the edge of the galaxy will be considered to be at the last observational velocity curve datapoint ($r_{d}$). In the case of $M_{DM}$, the galactic-distance-corrected distance is used. Using the insights from eqs. 13 and 16, by considering the fitted K parameter for the galaxy, it is possible to estimate the apparent asymptotic mass of the galaxy by setting the acceleration inferred from the edge of the galaxy to equal to the vacuum solution:
@@
In the case of $K=0$, eq. 65 should approximately yield the gravitating mass required for the observed velocity for the galaxy and is expected from the total mass inside $r_{d}$ with DM included. It can also be used to estimate the mass of the stellar disk (SD) and gas (HI) in the case of spherical symmetry, by using their respective velocities in the equation (and factoring in our fitted mass-luminosity ratio). Therefore, we can call $M_{aa,K=0}$ as apparent total asymptotic mass of the galaxy. The results are shown in Table 1:
Galaxy $M_{SD}$ $M_{HI}$ $M_{BM}$ $M_{DM}$ $M_{BM+DM}$ $M_{aa,K=0}$ $M_{aa}$
NGC 3198 $34.1$ $23.4$ $57.5$ $147.5$ $205.0$ $227.5$ $927.8$
Table 1. The mass components and estimates for DC14 DM mass, all values are in the units of $10^{9}M_{\odot}$. $M_{SD}$ is the mass of the stellar components, $M_{HI}$ is the mass of neutral hydrogen, $M_{BM}$ is the total mass of $M_{SD}$ and $M_{HI}$, $M_{DM}$ is the mass of the DC14 model DM halo up to $r_{d}$, $M_{BM+DM}$ is the total mass, including baryonic and DM components, and $M_{aa}$ is the apparent asymptotic mass from the thermodynamic theory.
We can see, that the total baryonic and dark matter mass of NGC 3198, $M_{BM+DM}$, is smaller than the apparent total mass of TG, but the apparent asymptotic mass with the galactic $K$ value, $M_{aa}$, while differing from the DC14 DM mass, up to $r_{d}$ results in the same gravitational effect. The characteristic $R$ spatial distance of the crossover, as pictured on Fig. 4 and detailed in [24] can be obtained as
(66) $R=KGM_{aa,K=0},$
@@
Using the $M_{aa,K=0}=227.5\cdot 10^{9}M_{\odot}$ mass for this scale, the expected characteristic crossover distance of this galaxy is $R\approx 33.3$ kpc. That is a crude estimate for distance where the flat velocity curve is expected to start declining and tending to a Newtonian one, which is determined by the vacuum solution of the Poisson equation with the apparent total asymptotic mass.
@@ 3.5. Conclusions
In this work, the analytical solutions of the nonlinear stationary field equation of TG were presented for the gravitational field in the cases of vacuum and constant matter density. Using the staggered grid-type discretisation, a numerical relaxation method was developed to solve the dissipative field equation for arbitrary spherically symmetric mass distribution. The stationary solution was obtained in the limit of the relaxation. The rotational velocity curve was calculated for galaxy NGC 3198 from the SPARC dataset, with only the baryonic matter contributing to the matter density source term [34]. The $K$ parameter in the theory investigated by this work arises from the coupling between the gravitational and mechanical thermodynamic forces and fluxes and, thus, naturally varies across the galaxies, depending on their smaller-scale dynamics and characteristics.
Based on the developed numerical method, a next step is to extend the analysis to other galaxies of the SPARC sample, with a detailed comparison of the alternative models, in particular for Fuzzy DM hydro solutions, similarly to [44] and [6]. It is also natural to investigate thermodynamic gravity on the scale of galaxy clusters. It may be the case that on much larger scales, the theory has effects similar to Newtonian or general relativistic gravity, except the apparent mass is larger due to the thermodynamic coupling, with a crossover between the regimes [24].
@@
For simplicity and analytical comparisons, the numerical method worked with a spherically symmetric model of mass distribution, although the baryonic mass of the non-elliptical galaxies usually is described by exponential disks with different scale heights in the axial $z$ and radial directions. However, the solution of nonlinear field equations for nonspherical mass distributions can be mathemathically demanding [48] and certain physical considerations should be taken into account [49, 50, 51].
@@ Acknowledgements
We would like to thank Professor W. J. G. de Blok for providing us with the THINGS data [52]. We would also like to thank S. Abe and R. Trasarti-Battistoni for providing insightful remarks regarding the theory. We would like to thank the anonymous referees for their constructive and helpful feedback.
@@
The data for the galaxy NGC 3198 used in this paper is based on the SPARC sample, [34].